OpenStudy (anonymous):

state how many imaginary and real zeros the function has. x^4-8x^3+17x^2-8x+16

OpenStudy (anonymous):

It's either 0 imaginary: 4 real or 2 imaginary: 2real

OpenStudy (anonymous):

\[x^4-8 x^3+17 x^2-8 x+16=(x-4)^2 \left(x^2+1\right) \]

OpenStudy (anonymous):

so its all real?

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

is it 1 real and 3 imaginary?

OpenStudy (anonymous):

someone please help

OpenStudy (lopezking1):

Do you have nay choices to go by just wondering

ganeshie8 (ganeshie8):

we can use Descartes rules of signs

OpenStudy (anonymous):

yes. 0 real: 4 imaginary 1 real: 3 imaginary 2 real: 2 imaginary 4 real : 0 imaginary

ganeshie8 (ganeshie8):

x^4-8x^3+17x^2-8x+16

OpenStudy (anonymous):

x=4 x=-i x=i

OpenStudy (anonymous):

nevermind guys i got it

ganeshie8 (ganeshie8):

great ! you can strike off the first option cuz 3 imaginary is never possible for any polynomial, cuz imaginary roots come in pairs

ganeshie8 (ganeshie8):

you need to figure out the answer from last two... which u already figured out it seems :)

OpenStudy (anonymous):

about 4

OpenStudy (anonymous):

lol idk the ratio tho :P